OpenStudy (anonymous):

for the given angle find the coterminal angle such that -360

OpenStudy (loser66):

I don't understand the question, your angle is a, b, c, d and which of them has the same terminal with 35? right?

OpenStudy (anonymous):

the ones under 35 are the answers

OpenStudy (loser66):

nope, still not get, @zepdrix

OpenStudy (loser66):

@amistre64

zepdrix (zepdrix):

It should be one of these two options. \[\large 35^o+360^o=395^o\]\[\large 35^o-360^o=-325^o\] Which one falls within the given interval? \(\large -360^o\) to \(\large 360^o\) ?

OpenStudy (anonymous):

b